OpenStudy (anonymous):

The table below shows four systems of equations: System 1 System 2 System 3 System 4 4x − 5y = 2 3x − y = 8 4x − 5y = 2 10x − 7y = 18 4x − 5y = 2 3x − 8y = 4 4x − 5y = 2 10x + 3y = 15 Which pair of systems will have the same solution? System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to two times the second equation in system 1. System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation

OpenStudy (anonymous):

Which pair of systems will have the same solution? System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to two times the second equation in system 1. System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation in system 2. System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to three times the second equation in system 1. System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to three times the second equation in system 2.

OpenStudy (anonymous):

@hoblos

OpenStudy (hoblos):

two systems have the same solution if the the equations of the second system are a combination of the equations of the first so you have to check each of the cases given, the first statement: second equation in system 2 is obtained by adding the first equation in system 1 to two times the second equation in system 1 4x − 5y = 2 + 2(3x − y = 8) --------------- = 10x -7y = 18 so the first choice is true

OpenStudy (anonymous):

thanks :D could you please help me with one more?

OpenStudy (hoblos):

post the question and i'll try